OpenStudy (phagey):

1.Write an inequality for all real numbers that are at least -1 and at most 3 2. A 130lb women and her 68lbs nephew step into the elevator and the women holds the door for a maintenance worker to wheel a 115lbs dolly with boxes stacked on top. The maintenance worker weighs 208lbs if the elevator can hold at most 100lbs write an inequality to describe how much additional weight the elevator can safely carry

OpenStudy (agent47):

\[-1 \le x \le 3\]

OpenStudy (agent47):

The second question is wrong Weight inside = 130 + 68 + 208 + 115 lbs. The elevator can carry (capacity - weight inside) more pounds.